# Emergence of non-Markovian Decoherent Histories in Integrable Environment: A “Tape Recorder” Model for Local Quantum Observables

### v2

@@ I Introduction
The paper is structured as follows. Section II provides preliminaries on decoherent histories and introduces the tape recorder model. Section III describes the physical system and its coarse-grained description. In Section IV, we introduce a numerical method based on out-of-time-order correlators (OTOCs) to identify the irreversibly decoupled environmental modes. Section V presents the computation of the decoherence functional. In Section VI, we demonstrate that our decoherent histories represent a stochastic unravelling of the exact quantum dynamics of the open system, and we present an efficient Monte Carlo simulation technique based on this insight. Finally, Sections VII and VIII summarize our results and present conclusions.
@@ II.2 Tape-recorder model as a physical mechanism
We emphasize that this division is approximate. Just as the head formally interacts nonlocally with the entire tape, decoherent histories only approximately satisfy the consistency conditions. The significance threshold enables this practical classification.
@@ III.1 Local interaction quench
$\hat{a}_{0}^{\dagger}\left(t\right)=\sum_{k=1}^{\infty}\alpha_{k}\left(t\right)\hat{a}_{k}^{\dagger},$ (17)
@@
As an example, in Fig. 2 (a) we provide a numerically exact calculation of chain site occupations for the case when the open system is a driven qubit: $\hat{H}_{s}\left(t\right)=\hat{\sigma}_{+}\hat{\sigma}_{-}+\hat{\sigma}_{x}\,f\cos t$ and the environment is bosonic. Here $f=0.1$ and other parameters of model eq. (13, 14) are: $g=0.05$, $\hat{V}_{s}=\hat{\sigma}_{-}$, $\varepsilon_{j}\equiv 1$, $h_{j}\equiv 0.05$. In Fig. 2 (b) we illustrate the light-cone-like spread of $\hat{a}_{0}^{\dagger}\left(t\right)$ by plotting $\left|\alpha_{k}\left(t\right)\right|$ as a function of chain site number and time with maximum Lieb-Robinson velocity $v_{L}=2\,h_{j}$.
@@
To realize this four-region coarse-graining, we transform the original chain modes $\hat{a}_{1}^{\dagger},\hat{a}_{2}^{\dagger},\ldots$ into new modes corresponding to the tape recorder domains. In this way, the environment is mapped onto a “moving tape”. This is achieved by a time-dependent Bogoliubov transformation, which naturally generates wave-packet–like localized modes:
$\hat{\kappa}_{p}^{\dagger}(t)=\sum_{q=1}^{\infty}U_{pq}(t)\hat{a}_{q}^{\dagger}$ (20)
with $\left[UU^{\dagger}\right]_{pq}=\delta_{pq}$. We refer these modes as follows (i) $\kappa_{p}^{\rm edge}$, (ii) $\kappa_{p}^{\rm in}$, (iii) $\kappa_{p}^{\rm rel}$, (iv) $\kappa_{p}^{\rm out}$ (Fig. 3).
@@ IV Computational Framework
In this section we will explicitly derive environmental degrees of freedom, which carry significant information about open system, together with the corresponding time moments. We show that the sequence of projections onto its subspace after escape times satisfy the approximate decoherence condition with increasing accuracy for decreasing significance threshold.
@@ IV.1 Irrelevant subspace for local interaction quench: edges of tape
It can be expressed in terms of $\left|\kappa\right\rangle$ :
$g_{+}\left(\kappa,T\right)=\left\langle\kappa\right|\hat{\rho}_{+}\left(T\right)\left|\kappa\right\rangle-a_{\rm cut}<0$ (23)
where we introduce the retarded light-cone density matrix:
@@
with $\alpha\left(\tau\right)$ being the solution of the operator spread equation (1819).
@@
$\textrm{if}\,\,g_{+}\left(\kappa,t\right)<0\,\,\textrm{then}\,\,g_{+}\left(\kappa,t^{\prime}\right)<0\,\,\textrm{for all}\,\,t^{\prime}<t,$ (25)
@@
$\sqrt{F(\kappa,t)}=\left|\langle\Psi_{\perp}(t)|\Psi(t)\rangle\right|$ (29)
@@
$I\left(\kappa,t\right)=1-\sqrt{F\left(\kappa,t\right)}$ (30)
yields a measure of the significance of $\kappa$ during $\left[0,t\right]$ and converges to zero for modes outside light-cone interior, correspondind the edges of tape.
@@
If we include $m\left(t\right)=Bt+\delta$ light cone interior normal modes, the resulting infidelity is:
$I\left(\kappa,t\right)\lesssim\left[g\left\|\hat{V}_{s}\right\|\right]^{2}\mathcal{I}_{1}^{+}\left(\infty\right)e^{-\gamma\delta}$ (32)
i.e. exponential convergence with respect to the number $\delta$ of additional “guard” modes. Here $B$, $\gamma$ is some positive constant (see Appendix A). Equivalently, the choice $\delta$ corresponds to a relative cutoff $r_{\rm cut}=e^{-\gamma\delta}$.
@@ IV.2 Basis of a minimal light cone normal modes
$t_{\rm in}\left(\kappa\right)=\int_{0}^{T}d\tau\,\theta\left[\,-\,g_{+}\left(\kappa,\tau\right)\right]$ (33)
@@
Figure 4: The number of arrived modes $m_{+}\left(t\right)$ vs time $t$ on different scales. $m_{+}\left(t\right)$ is defined as the number of modes $\kappa$ such that $t_{\rm in}\left(\kappa\right)\leq t$. The parameters used are: $a_{\rm cut}=10^{-5},\,\varepsilon_{j}\equiv 1$, $h_{j}\equiv 0.05$. The chain sites couple one-by-one at an asymptotically constant rate (blue curve). At the same time, the light cone interior normal modes couple almost instantly at $t=0$ (green curve).
Following the tape recorder analogy (Sec. II.2), we retain only the modes that carry statistically significant records (i.e., the light-cone interior normal modes) and that also couple sequentially, as in the chain sites. To optimize the localization of independent degrees of freedom, we rotate the basis:
$\hat{\bar{\kappa}}_{p}^{\dagger}=\sum_{q=1}^{m(T)}U_{pq}\hat{\widetilde{\kappa}}_{q}^{\dagger},\quad p=1,\ldots,m(T),$ (34)
@@
and $\hat{\kappa}_{p}^{\rm in\dagger}$ are such modes, whose arrival times $t_{\rm in}\left(\kappa_{p}^{\rm in}\right)$ are delayed as much as possible. These modes contain significant information about the open system, the decoherent history is recorded in them (ii).
Thus, on a finite time interval $\left[0,T\right]$ the full system dynamics (13) can be approximated by effective Hamiltonian $H_{+\rm min}(t)$ only with $m_{+}(t)$ relevant modes (Appendix B):
@@
Figure 5: Square-root infidelity between the state $\left|\Psi\left(T\right)\right\rangle^{\prime}$ obtained for $\hat{H}_{+\rm min}\left(t\right)$ eq. (36), and the state $\left|\Psi\left(T\right)\right\rangle$ containing all modes $\hat{\kappa}_{k}^{\rm in}$. The nonstationary Schrodinger equation was solved numerically in a Fock space which was truncated at a maximal number of environment’s quanta $n_{\rm cut}=4$ for $T=100$. The driven qubit $\hat{H}_{s}\left(t\right)=\hat{\sigma}_{+}\hat{\sigma}_{-}+\hat{\sigma}_{x}0.1\cos t$ was taken as an open system, $g=0.1$, $\hat{V}_{s}=\hat{\sigma}_{-}$, $\varepsilon_{j}\equiv 1$, $h_{j}\equiv 0.05$. The semiinfinite chain eq. (14) was truncated at 3030 sites.
@@ IV.3 Minimal backward light cone with stable records
$\int\limits_{t}^{T}d\tau\,\,\mathcal{C}\left(\kappa,\tau\right)=\left\langle\kappa\right|\hat{\rho}_{-}\left(t\right)\left|\kappa\right\rangle$
where $\hat{\rho}_{-}\left(t\right)=\int\limits_{t}^{T}d\tau\left|\alpha\left(\tau\right)\right\rangle\left\langle\alpha\left(\tau\right)\right|$ with $\alpha\left(\tau\right)$ being the solution of eq. (18-19).
@@
$\left[\rho_{-}^{m_{+}\left(t\right)}\left(t\right)\right]_{pq}=\left\langle\kappa_{p}^{\rm in}\left|\hat{\rho}_{-}\left(t\right)\right|\kappa_{q}^{\rm in}\right\rangle\,\,\textrm{for}\,\,p,q=\overline{1,m_{+}\left(t\right)}\,\,.$
The contribution of such modes to the system’s future dynamics becomes negligible after time $t$.
@@ V Decoherence overlap
$\begin{split}\hat{H}\left(t\right)\approx\hat{H}_{\rm eff}(t)=\hat{H}_{s}+g\,\hat{V}_{s}^{\dagger}\sum_{k=1}^{r(t)}\chi_{k}^{*}\left(t\right)\hat{\kappa}_{k}^{\rm rel}\,+\\ +\,g\,\hat{V}_{s}\sum_{k=1}^{r\left(t\right)}\chi_{k}\left(t\right)\hat{\kappa}_{k}^{\rm rel}-\sum_{k=1}^{r(t)}\sum_{l=1}^{r\left(t\right)}\xi_{kl}\left(t\right)\hat{\kappa}_{k}^{\rm rel}{}^{\dagger}\hat{\kappa}_{l}^{\rm rel}\end{split}$ (41)
where $\chi_{k}(t)$ are the time-dependent coupling amplitudes, $\xi_{kl}$ gradually implement the rotation and together with $\hat{\kappa}_{k}$ are defined by the non-stationary Bogoliubov transformation Pol2022 .
Figure 6: Average decoherence overlap (7) for different significance threshold $r_{\rm cut}$. Here no truncation to the relevant modes was applied; otherwise the overlaps would vanish exactly after each projection.
@@
$\left[\hat{H}_{\rm int}(t),\hat{P}_{{\kappa^{\rm out}};\,n_{\kappa^{\rm out}}}\left(\,t(\kappa^{\rm out})\,\right)\right]\approx 0$ (42)
for $t>t(\kappa^{\rm out})$ up to an accuracy controlled by significance threshold $r_{\rm cut}$. For the projectors we use occupation number eigensubspace:
$\hat{n}_{p}^{\rm out}\hat{P}_{{\kappa_{p}^{\rm out}};\,n_{p}^{\rm out}}=n_{p}^{\rm out}\hat{P}_{{\kappa_{p}^{\rm out}};\,n_{p}^{\rm out}}$
@@
with pure initial condition (15). Here $\hat{U}(t_{i},t_{j})=\exp(-i\hat{H}(t_{i}-t_{j}))$ is full evolution.
@@
up to the accuracy controlled by the significance threshold, demonstrating approximate diagonality.
@@
Fig. 6 shows the average decoherence overlap (44) computed over 20 sampled histories, while Fig. 7 illustrates its time dependence for different significance thresholds $r_{\rm cut}$. These results confirm that the identified modes are indeed irreversibly decoupled and that projections onto their subspaces yield decoherent histories.
Figure 7: Average decoherence overlap for different 35 histories depending on time. The number of modes inside minimal light cone computed with different significance threshold $r_{\rm cut}$. The Fock space was truncated at a maximal number quanta $n_{\rm cut}=5$ for $T=150$. The driven qubit $\hat{H}_{s}\left(t\right)=\hat{\sigma}_{+}\hat{\sigma}_{-}+\hat{\sigma}_{x}\,0.1\,\cos t$ was taken as an open system, $g=0.1$, $\hat{V}_{s}=\hat{\sigma}_{-}$, $\varepsilon_{j}\equiv 1$, $h_{j}\equiv 0.05$. Here no truncation to the relevant modes was applied; otherwise the overlaps would vanish exactly after each projection.

### v3

@@ I Introduction
The paper is structured as follows. Section II provides preliminaries on decoherent histories and introduces the tape recorder model. Section III describes the physical system and its coarse-grained description. In Section IV, we introduce a numerical method based on out-of-time-order correlators (OTOCs) to identify the irreversibly decoupled environmental modes. Section V presents the computation of the decoherence functional and discusses the physical interpretation of the significance threshold. In Section VI, we demonstrate that our decoherent histories represent a stochastic unravelling of the exact quantum dynamics of the open system, and we present an efficient Monte Carlo simulation technique based on this insight. Finally, Sections VII and VIII summarize our results and present conclusions.
@@ II.2 Tape-recorder model as a physical mechanism
We emphasize that this division is approximate. Just as the head formally interacts nonlocally with the entire tape, decoherent histories only approximately satisfy the consistency conditions. The significance threshold enables this practical classification, and is not a free parameter. The finite resolution of the read-write head admits a quantum analogue in the form of the experimental infidelity, which reflects unavoidable imperfections in state preparation, evolution, or measurement. Once residual coherences fall below this level, quantum superpositions become operationally indistinguishable from classical mixtures. The significance threshold therefore is determined directly by the experimentally achievable infidelity.
@@ III.1 Local interaction quench
$\hat{a}_{0}^{\dagger}(t)=\sum_{k=0}^{\infty}\alpha_{k}(t)\hat{a}_{k}^{\dagger}$ (17)
@@
[…]
Figure 3: Time dependence for the trace distance $T(\rho_{1},\rho_{2};t)=\dfrac{1}{2}\sqrt{(\rho_{1}(t)-\rho_{2}(t))^{2}}$ for different parameters of model (13, 14). For coupling $g=0.05$ and a qubit frequency at the center of the environmental band ($\Omega=1$), weak oscillations are observed, with a BLP measure of $0.13$. Shifting the qubit frequency toward the band edge ($\Omega=0.9$) makes the oscillations more visible, while the BLP value is $0.05$. For stronger coupling ($g=0.1$), large-amplitude oscillations appear and the BLP measure increases to $1.19$, indicating a strongly non-Markovian regime. As the BLP is a scalar quantity, the full time dependence of the trace distance provides a more complete characterization of the dynamics.
As an example, in Fig. 2 (a) we provide a numerically exact calculation of chain site occupations for the case when the open system is a driven qubit: $\hat{H}_{s}\left(t\right)=\Omega\,\hat{\sigma}_{+}\hat{\sigma}_{-}+\hat{\sigma}_{x}\,f\cos t$ and the environment is bosonic. Here $f=0.1$ and other parameters of model eq. (13, 14) are: $\Omega=1$, $g=0.05$, $\hat{V}_{s}=\hat{\sigma}_{-}$, $\varepsilon_{j}\equiv 1$, $h_{j}\equiv 0.05$. In Fig. 2 (b) we illustrate the light-cone-like spread of $\hat{a}_{0}^{\dagger}\left(t\right)$ by plotting $\left|\alpha_{k}\left(t\right)\right|$ as a function of chain site number and time with maximum Lieb-Robinson velocity $v_{L}=2\,h_{j}$.
The dynamics in this model is generally non-Markovian. To quantify non-Markovianity we evaluate the time-dependent trace distance $T(\rho_{1},\rho_{2};t)=\dfrac{1}{2}\sqrt{(\rho_{1}(t)-\rho_{2}(t))^{2}}$ for two orthogonal initial qubit states $|\rho_{1}(0)=\ket{\uparrow}\bra{\uparrow}$ and $|\rho_{2}(0)=\ket{\downarrow}\bra{\downarrow}$. The trace distance exhibits clear non-monotonic behavior, indicating non-Markovian dynamics and information backflow from the environment to the system Breuer2009 (see Fig. 3).
@@
To realize this four-region coarse-graining, we transform the original chain modes $\hat{a}_{0}^{\dagger},\hat{a}_{1}^{\dagger},\ldots$ into new modes corresponding to the tape recorder domains. In this way, the environment is mapped onto a “moving tape”. This is achieved by a time-dependent Bogoliubov transformation, which naturally generates wave-packet–like localized modes:
$\hat{\kappa}_{p}^{\dagger}(t)=\sum_{q=0}^{\infty}U_{pq}(t)\hat{a}_{q}^{\dagger}$ (21)
with $\left[UU^{\dagger}\right]_{pq}=\delta_{pq}$. We introduce the corresponding one particle quantum state:
[…]
@@ IV Computational Framework
In this section we will explicitly derive environmental degrees of freedom, which carry significant information about open system, together with the corresponding time moments. We show that the sequence of projections onto its occupation number eigensubspace after escape times satisfy the approximate decoherence condition with increasing accuracy for decreasing significance threshold.
@@ IV.1 Irrelevant subspace for local interaction quench: edges of tape
This condition can be conveniently rewritten using the state $|\kappa\rangle$. Indeed, by introducing the retarded light-cone density matrix:
@@
with $\alpha\left(\tau\right)$ being the solution of the operator spread equations (1819) and using the relation:
$\begin{gathered}\left[\hat{a}_{0}\left(\tau\right),\hat{\kappa}^{\dagger}\right]=\left[\sum_{q=0}^{\infty}\alpha_{q}^{*}(\tau)\hat{a}_{q},\sum_{p=0}^{\infty}\kappa_{p}\hat{a}_{p}^{\dagger}\right]=\\ =\sum_{p=0}^{\infty}\alpha_{p}^{*}(\tau)\kappa_{p}=\langle\alpha(\tau)|\kappa\rangle\end{gathered}$ (26)
which follows from (17) and (21), the average intensity takes the form:
[…]
Consequently, the condition (24) becomes:
$g_{+}\left(\kappa,T\right)=\left\langle\kappa\right|\hat{\rho}_{+}\left(T\right)\left|\kappa\right\rangle-a_{\rm cut}<0\,.$ (27)
@@
$\textrm{if}\,\,g_{+}\left(\kappa,t\right)<0\,\,\textrm{then}\,\,g_{+}\left(\kappa,t^{\prime}\right)<0\,\,\textrm{for all}\,\,t^{\prime}<t$ (28)
@@
where $\mathcal{I}_{1}^{+}\left(t\right)\equiv\mathcal{I}^{+}\left(\widetilde{\kappa}_{1},t\right)$. The threshold $a_{\rm cut}$ can be specified by choosing a relative threshold $r_{\rm cut}$ (e.g. $r_{\rm cut}=10^{-4}$):
$a_{\rm cut}=r_{\rm cut}\,\mathcal{I}_{1}^{+}\left(t\right)\,.$ (31)
@@
$\sqrt{F(t)}=\left|\langle\Psi_{\perp}(t)|\Psi(t)\rangle\right|$ (33)
@@
$I\left(t\right)=1-\sqrt{F\left(t\right)}$ (34)
yields a measure of the significance of $\kappa$ during $\left[0,t\right]$ and converges to zero for modes outside light-cone interior, corresponding the edges of tape.
@@
If we consider the analogy with the Lieb–Robinson light cone in mode space and include $m\left(t\right)=Bt+\delta$ light cone interior normal modes, the resulting infidelity is:
$I\left(t\right)\lesssim\left[g\left\|\hat{V}_{s}\right\|\right]^{2}\mathcal{I}_{1}^{+}\left(\infty\right)e^{-\gamma\delta}$ (36)
and does not depend on time, providing uniform exponential convergence with respect to the number $\delta$ of additional “guard” modes. Here $B$, $\gamma$ is some positive constant (see Appendix A). Equivalently, the choice $\delta$ corresponds to a relative cutoff $r_{\rm cut}=e^{-\gamma\delta}$.
@@ IV.2 Basis of a minimal light cone normal modes
$t_{\rm in}\left(\kappa\right)=\int\limits_{0}^{T}d\tau\,\theta\left[\,-\,g_{+}\left(\kappa,\tau\right)\right]$ (37)
@@
Figure 5: The number of arrived modes $m_{+}(t)$ as a function of time $t$ is shown for different choices of basis. Here $m_{+}(t)$ is defined as the number of modes $\kappa$ such that $t_{\rm in}(\kappa)\leq t$. The parameters used are: $a_{\rm cut}=10^{-5}$, $\varepsilon_{j}\equiv 1$, $h_{j}\equiv 0.05$. The behavior of $m_{+}(t)$ is strongly basis dependent. In the chain basis, the modes couple one-by-one at an asymptotically constant rate (blue curve). In contrast, in the normal-mode basis corresponding to the light-cone interior normal modes, the modes couple to the system almost instantaneously at $t\approx 0$ (green curve).
@@
Following the tape recorder analogy (Sec. II.2), we retain only the modes that carry statistically significant records (i.e., the light-cone interior normal modes) and that also couple sequentially, as in the chain sites. To optimize the localization of independent degrees of freedom, we rotate the normal mode basis:
$\hat{\bar{\kappa}}_{p}^{\dagger}=\sum_{q=1}^{m(T)}U_{pq}\hat{\widetilde{\kappa}}_{q}^{\dagger},\quad p,q=1,\ldots,m(T),$ (38)
@@
and $\hat{\kappa}_{p}^{\rm in\dagger}$ are such modes, whose arrival times $t_{\rm in}\left(\kappa_{p}^{\rm in}\right)$ are delayed as much as possible. Details of the construction are given in the Appendix C and Pol2022 . These modes contain significant information about the open system, the decoherent history is recorded in them (ii).
Thus, on a finite time interval $\left[0,T\right]$ the full system dynamics (13) can be approximated by effective Hamiltonian $H_{+\rm min}(t)$ in the frame of the minimal light cone only with $m_{+}(t)$ light cone modes (Appendix B):
@@
Figure 6: Square-root infidelity between the state $\left|\Psi\left(T\right)\right\rangle^{\prime}$ obtained for $\hat{H}_{+\rm min}\left(t\right)$ eq. (40), and the state $\left|\Psi\left(T\right)\right\rangle$ containing all modes $\hat{\kappa}_{k}^{\rm in}$. The nonstationary Schrodinger equation was solved numerically in a Fock space which was truncated at a maximal number of environment’s quanta $n_{\rm cut}=4$ for $T=100$. The driven qubit $\hat{H}_{s}\left(t\right)=\hat{\sigma}_{+}\hat{\sigma}_{-}+\hat{\sigma}_{x}0.1\cos t$ was taken as an open system, $g=0.1$, $\hat{V}_{s}=\hat{\sigma}_{-}$, $\varepsilon_{j}\equiv 1$, $h_{j}\equiv 0.05$. The semiinfinite chain eq. (14) was truncated at $30$ sites.
@@ IV.3 Minimal backward light cone with stable records
$\int\limits_{t}^{T}d\tau\,\,\mathcal{C}\left(\kappa,\tau\right)=\left\langle\kappa\right|\hat{\rho}_{-}\left(t\right)\left|\kappa\right\rangle$ (42)
where $\hat{\rho}_{-}\left(t\right)=\int\limits_{t}^{T}d\tau\left|\alpha\left(\tau\right)\right\rangle\left\langle\alpha\left(\tau\right)\right|$ with $\alpha\left(\tau\right)$ being the solution of eqs. (18-19).
@@
$\left[\rho_{-}^{m_{+}\left(t\right)}\left(t\right)\right]_{pq}=\left\langle\kappa_{p}^{\rm in}\left|\hat{\rho}_{-}\left(t\right)\right|\kappa_{q}^{\rm in}\right\rangle$ (44)
for $p,q=1,\ldots,m_{+}\left(t\right)$. The contribution of such modes to the system’s future dynamics becomes negligible after time $t$.
@@
Figure 7: The plot shows the average decoherence overlap (7) over 50 sampled histories as a function of time, for different values of the significance threshold $r_{\rm cut}$. The number of modes inside minimal light cone computed with different significance threshold $r_{\rm cut}$. The Fock space was truncated at a maximal number quanta $n_{\rm cut}=5$ for $T=150$ and $T=200$. The driven qubit $\hat{H}_{s}\left(t\right)=\hat{\sigma}_{+}\hat{\sigma}_{-}+\hat{\sigma}_{x}\,0.1\,\cos t$ was taken as an open system, $g=0.05$, $\hat{V}_{s}=\hat{\sigma}_{-}$, $\varepsilon_{j}\equiv 1$, $h_{j}\equiv 0.05$. Here no truncation to the relevant modes was applied; otherwise the overlaps would vanish exactly after each projection.
@@ V Decoherence overlap
$\begin{split}\hat{H}\left(t\right)\approx\hat{H}_{\rm eff}(t)=\hat{H}_{s}+g\,\hat{V}_{s}^{\dagger}\sum_{k=1}^{r(t)}\chi_{k}^{*}\left(t\right)\hat{\kappa}_{k}^{\rm rel}\,+\\ +\,g\,\hat{V}_{s}\sum_{k=1}^{r(t)}\chi_{k}\left(t\right)\hat{\kappa}_{k}^{\rm rel}{}^{\dagger}-\sum_{k=1}^{r(t)}\sum_{l=1}^{r(t)}\xi_{kl}\left(t\right)\hat{\kappa}_{k}^{\rm rel}{}^{\dagger}\hat{\kappa}_{l}^{\rm rel}\end{split}$ (47)
where $\chi_{k}(t)$ are the time-dependent coupling amplitudes, $\xi_{kl}(t)$ gradually implement the rotation and together with $\hat{\kappa}_{k}$ are defined by the non-stationary Bogoliubov transformation (see Appendix D, Pol2022 ).
@@
$\left[\hat{H}_{\rm int}(t),\hat{P}_{{\kappa^{\rm out}};\,n_{\kappa^{\rm out}}}\right]=O\left(\sqrt{r_{\rm cut}}\right)\approx 0$ (48)
for $t>t_{\rm out}(\kappa^{\rm out})$ up to an accuracy controlled by significance threshold $r_{\rm cut}$ (see Appendix E). For the projectors we use occupation number eigensubspace:
$\hat{n}_{p}^{\rm out}\hat{P}_{{\kappa_{p}^{\rm out}};\,n_{p}^{\rm out}}=n_{p}^{\rm out}\hat{P}_{{\kappa_{p}^{\rm out}};\,n_{p}^{\rm out}}$ (49)
@@
Figure 8: The time dependence of the mean plateau value of the decoherence functional. Here no truncation to the relevant modes was applied; otherwise the overlaps would vanish exactly after each projection.
@@
with pure initial condition (15). Here $\hat{U}(t_{i},t_{j})$ is full time evolution.
@@
up to the accuracy controlled by the significance threshold, demonstrating approximate diagonality (see Appendix E).
@@
The Fig. 7 illustrates the average decoherence overlap (7) time dependence for different significance thresholds $r_{\rm cut}$ computed over 50 sampled histories. The Fig. 8 shows the average decoherence overlap plateau depending on $r_{\rm cut}$, computed over 50 sampled histories. These results confirm that the identified modes are indeed irreversibly decoupled and that projections onto their subspaces yield decoherent histories.
Importantly, irreversibly decoupled modes appear at characteristic times that remain essentially unchanged upon increasing the total evolution time $T$ (see Appendix D, Figs. 14,15).This ensures that the decoherence overlap remains stable for finite-time simulations.
V.1 Physical meaning of significance threshold $r_{\rm cut}$
As demonstrated by the infidelity of many-body joint state of system and environment $I(t)$ (34, 36), the contribution of the discarded modes to the dynamics is proportional to $r_{\rm cut}$, and can be made arbitrarily small by decreasing the threshold, ensuring that observables remain accurately reproduced.
Figure 9: Time evolution of $\langle\sigma_{z}\rangle$. The irreversibly decoupled modes are traced out using the quantum jump Monte Carlo simulation, providing an efficient unraveling of the full wave function into a sampled ensemble of 300 decoherent histories. The value of the root mean square deviation (RMS) is shown in the figure corresponding to the range of times from 0 to 125 and decrease with $r_{\rm cut}$.
Conceptually, $r_{\rm cut}$ has an operational meaning: it is determined by the smallest coherences that can be resolved in an experiment. Any measurement setup has finite time resolution, bandwidth, and signal-to-noise ratio, while environmental effects such as thermal fluctuations further limit coherence detection Clerk2010 , setting a lower bound on the experimentally achievable infidelity. Residual coherences between different decoherent histories then lie below this resolution and are operationally unobservable. This defines a physically meaningful distinguishability scale, which naturally identifies the appropriate range of $r_{\rm cut}$.
Concretely, $r_{\rm cut}$ should be chosen from infidelity bound (34):
$r_{\rm cut}\propto a_{\rm cut}=\dfrac{I_{\rm experimental}}{\left(g\|\hat{V}_{s}\|\right)^{2}}$ (52)
where $I_{\rm experimental}$ is the experimentally achievable infidelity.
Importantly, the precise numerical value of $r_{\rm cut}$ is not critical. Observables are robust as shown in Fig. 9, and the escape time $t^{\rm out}$ depends only logarithmically on $r_{\rm cut}$ (see Fig. 10), so variations of the threshold within the experimentally admissible range produce only minor changes. This demonstrates that the emergence of classical records is insensitive to the exact choice of $r_{\rm cut}$, reflecting the objective and operational nature of the classical limit.
Figure 10: The dependence of escape time $t^{\rm out}$ on $r_{\rm cut}$.